#1) A Submarine manufacturer is interested in continuing business of building mini-submarines bidding for the government to build mini-submarines. The company has built the first mini-submarine. The data for the first mini-submarine is as follows: Labor hours = 6000 Material cost = $200000 Labor cost = $50 per hour The company expects a Learning Rate of 80% in building these mini-submarines.

Compute total cost of producing the five mini-submarines.

Please show how it is solved clearly.
Answer #1

Time taken to produce nth submarine =Time taken to produce first submarine x ( n)b

where b is a constant which is given by b= log(learning rate) / log 2

For 80% learning  rate b = log(0.8) / log 2 = -0.3219

Time taken for producing second submarine =0.8x6000 =4800

Time taken for producing thrid submarine = 6000x (3)-0.3219 = 4212.75

Time taken for producing fourth submarine = 6000x (4)-0.3219 =3840.14

Time taken for producing fifth submarine = 6000x (5)-0.3219 =3574

Total hours taken by five submarines = 22426.89 =22427

Labour cost = 50x22427 = 1121350

Material cost =200000x5 =1000000 ( It is assumed that the learning cost is for labour only and the material will come at fixed price)

Total cost = 1000000+1121350 = 2121350